# The discrete periodic Pitman transform: invariances, braid relations, and Burke properties

### v2

@@ 1.3. Invariance of periodic polymer partition functions
$\mathcal{P}_{k}(\ldots,\mathbf{W}_{k-1},\mathbf{W}_{k},\mathbf{W}_{k+1},\mathbf{W}_{k+2},\ldots)=(\ldots,\mathbf{W}_{k-1},T(\mathbf{W}_{k+1},\mathbf{W}_{k}),D(\mathbf{W}_{k+1},\mathbf{W}_{k}),\mathbf{W}_{k+2},\ldots)$ (1.9)
Our first main result gives the invariance of the polymer partition functions under the iterated Pitman transform defined in (1.2).
@@
Let $N\in\mathbb{N}$, and let $\mathbf{W}=(\mathbf{W}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(W_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights. Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition $W_{k,j}=W_{k,\ell}$ whenever $j\equiv\ell\mod N$. If $k_{1},k_{2},\ldots,k_{m}\in\mathbb{Z}$ and the sets of initial points $U=\bigl{(}\bigl{(}a_{i},b_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ and end points $V=\bigl{(}(c_{i},d_{i})_{i\in\llbracket 1,m\rrbracket}\bigr{)}$ satisfy $(U,V)\in\Psi$ and, for all $i,j\in\llbracket 1,m\rrbracket$, $a_{i}\neq k_{j}+1$ and $c_{i}\neq k_{j}$, then,
$Z^{\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}}(V\mid U)=Z^{\mathbf{W}}(V\mid U).$ (1.10)
Remark 1.5.
See Figure 3 for a visual depiction of the invariance in Theorem 1.4. In Theorem 1.6 below, we show braid relations for the operators $\mathcal{P}_{k}$. Along with the fact that each $\mathcal{P}_{k}$ is an involution (see Lemma A.2), these relations imply that, if we have the equality of products of transpositions $(k_{m},k_{m}+1)(k_{m-1},k_{m-1}+1)\cdots(k_{1},k_{1}+1)=(k_{m}^{\prime},k_{m}^{\prime}+1)(k_{m-1}^{\prime},k_{m-1}^{\prime}+1)\cdots(k_{1}^{\prime},k_{1}^{\prime}+1)$, then $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}\mathbf{W}=\mathcal{P}_{k_{m}}^{\prime}\ldots\mathcal{P}_{k_{1}}^{\prime}\mathbf{W}$. Hence, we consider the action of $\mathcal{P}_{k_{m}}\ldots\mathcal{P}_{k_{1}}$ on $\mathbf{W}$ as the action of a general permutation, independent of its choice of decomposition. In this case, (1.11) gives an alternate condition for the set of endpoints whose partition function is preserved. An example is given in Figure 4 below. This connection is made precise in Lemma 4.2.
Figure 3. A depiction of the polymer invariance in Theorem 1.4. Each column is associated to a periodic vector of weights, $\mathbf{W}_{k}$. The horizontal lines are used to indicate a periodic environment for the polymer with $N=4$. In this example, we take $m=1$ and consider the action of the operator $\mathcal{P}_{3}$, which replaces the weights in columns $3$ and $4$ with $T(\mathbf{W}_{4},\mathbf{W}_{3})$ and $D(\mathbf{W}_{4},\mathbf{W}_{3})$, respectively. The partition function for each pair of points and the multi-path partition function for this collection of 33 pairs are both preserved under this transformation of the weights.
@@
Theorem 1.4 is proved in Section 2. It is the first of its kind for a polymer in a periodic environment. The invariance under Pitman transformations in other contexts has been widely studied. The concept can be traced back to queuing theory from the works [Web79, TW87], where it was shown that the law of the departures process from two $\cdot/M/1$ queues in series is invariant under swapping of the service rates of the two queues. In the polymer context, this type of result can be traced back to the work of Noumi and Yamada [NY04], who gave a description of the geometric Robinson-Schendsted-Knuth correspondence in terms of certain matrix product relations. Corwin [Cor21] used this to give an analogous result to our Theorem 1.4 for a full-line version of the discrete Pitman transform. The proof of Theorem 1.4 likewise uses the machinery from [NY04], but we instead work with $\mathbb{Z}\times\mathbb{Z}$ periodic matrices instead of finite-dimensional matrices.
@@
Theorem 1.4 limits to a zero-temperature analogue in Theorem 1.10. This zero-temperature limit is the periodic analogue of invariances shown in [DOV22, DNV22, DV24, BEM+25a]. The result in [DOV22] is particularly significant, as it was key in constructing the directed landscape, the central limiting object of the KPZ universality class.
@@
It would be interesting to see if Theorem 1.4 could be the starting point for constructing and proving convergence to the periodic directed landscape, but we note that the action of this transform is qualitatively different than that in [DOV22]. Specifically, in the full-line case, the action of the iterated Pitman transforms and independent collection of Brownian motions and transforms them to the Brownian melon–a set of nonintersecting paths whose geodesic structure is more amenable to asymptotic analysis. By contrast, in this periodic setting when we take the weights to be i.i.d. with the inverse-gamma distribution, the transformed weights are equal in distribution to the original weights. This is the Burke property shown in [CGS24, Proposition 5.12], and extended to an inhomogeneous setting in Proposition 4.1. As a corollary, we obtain a permutation invariance result for the multi-path inhomogeneous inverse-gamma polymer below in Theorem 1.7. This extends the previous result of [BEM+25a, Theorem 3.1] to a positive-temperature and multi-path setting. In this sense, the periodic Pitman transform is more analogous to a limiting case of the full-line Pitman transform, where the initial point of a random walk is sent to -$\infty$ and the output of the transform is recentered at 0 (see [DNV22, DV24]). This limiting full-line setting is where we see analogous Burke properties, as in [OY01, OY02].
1.4. Braid relations
We now describe our next result, which gives a braid relation for the Pitman operators defined in (1.9). The next theorem gives a braid relation for these operators, which is the same as that satisfied for the full-line Pitman transform [BBO05]. See also [DV24]. Combined with the fact that the map $\mathcal{P}_{k}$ is an involution (see Lemma A.2), the operators $\mathcal{P}_{k}$ generate a representation of the infinite symmetric group (note that the same braid relations are satisfied by the nearest-neighbor transpositions $(k,k+1)$).
@@ 1.4. Braid relations
For $k\in\mathbb{Z}$,
Theorem 1.6 is proved in Section 3. The key is the matrix relation in Corollary 2.5 below, which is also used to prove Theorem 1.4. When written out, the braid relation consists of three equalities ((3.1)-(3.3) below). The third of these equalities was previously proved in [CGS24, Proposition 5.13]. With that as our starting point, we iterate the matrix relation in Corollary 2.5 in two different ways to obtain and use the interpretation of the matrix product as the polymer partition function to obtain the other two equalities. Specifically, Lemma 3.1 is the key innovation that allows us to complete the proof.
1.5. Permutation invariance
In this section, we consider an inhomogeneous, periodic inverse-gamma polymer. Let $\lambda>0$ and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$. We take independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.15 below). Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$. We call a permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$ a finite permutation if $\sigma$ fixes all but finitely many coordinates. For a finite permutation $\sigma$, we let $\mathcal{U}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
@@ 1.5. Permutation invariance
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ satisfying $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$, define periodic weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}$ as above and extend them to $\mathbb{Z}\times\mathbb{Z}$. Then, for any finite permutation $\sigma:\mathbb{Z}\to\mathbb{Z}$, we have
$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma}\Bigr{)},$
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
The condition (1.11) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4 for an example.
Figure 4. An example of an element $(U,V)\in\mathcal{U}^{\sigma}$, where $U$ and $V$ each contain 33 elements. The gray horizontal lines are used to indicate a periodic environment with $N=4$. In this figure, the column parameters of columns $3,5$, and 66 (blue) are permuted, and the column parameters in columns $8$ and 99 (red) are permuted. The condition is that, for each of the three endpoints $(x,y)$ in $U$ (lower left), $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, while for each of the three endpoints $(x,y)$ in $V$ (upper right), $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$. The distributional equality in Theorem 1.7 holds jointly over all endpoint pairs $(U,V)\in\Psi$ satisfying this property.
We can allow for permuting the $\mathbf{b}$ parameters as well, with doubly periodic weights $(W_{(i,j)}\mid i\in\mathbb{Z}_{M},j\in\mathbb{Z}_{N})$ for $M,N\in\mathbb{N}$ when we restrict to points that lie inside the box $\llbracket 0,M\rrbracket\times\llbracket 0,N\rrbracket$. In this case, the model isn’t actually periodic at all; it is just a restriction of the full-space polymer to a finite grid. For finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, define the set $\mathcal{U}^{\sigma,\tau}$ to be the set of tuples $(U,V)$, where $U,V$ are each sets of endpoint pairs in $\mathbb{Z}\times\mathbb{Z}$ satisfying (1.11) as well as
[…]
The following is an extension of the recent result [BEM+25a, Theorem 3.1] to multi-paths and to the positive-temperature case. A sketch of the proof for the extension to multi-paths in zero temperature was previously discussed in [BEM+25a, Remark 4.3]. The technique follows the same outline as developed in [BEM+25a], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns. However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform. The algebra involved in this positive-temperature case is also different.
Theorem 1.8.
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}}$ satisfying $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}$ define independent weights $\mathbf{W}^{\mathbf{a},\mathbf{b}}=\bigl{(}W_{(i,j)}:i,j\in\mathbb{Z}\bigr{)}$ such that $W_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for all $i,j$. Then, for any finite permutations $\sigma,\tau:\mathbb{Z}\to\mathbb{Z}$, we have
$\Bigl{(}Z^{\mathbf{W}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)}\overset{d}{=}\Bigl{(}Z^{\mathbf{W}^{\sigma\mathbf{a},\tau\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{U}^{\sigma,\tau}\Bigr{)},$
@@
Theorems 1.7 and 1.8 are proved in Section 4. When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COSZ14, OSZ14, BC14]). A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [Dau22, Theorem 1.5], motivated by a shift-invariance result previously proved in Borodin, Gorin, and Wheeler [BGW22]. Connections between these results are discussed further in [BEM+25a, Section 3].
@@ 1.6. Zero-temperature limits
We also define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})$ $=\bigl{(}\widetilde{T}(\mathbf{W}_{2},\mathbf{W}_{1}),\widetilde{D}(\mathbf{W}_{2},\mathbf{W}_{1})\bigr{)},\quad\text{where}$ (1.13)
$\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
$\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{2,\ell+1}-X_{1,\ell}.$
The mapping $\widetilde{D}$ is similar to, but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07]. See [CGS24, Appendix F] for more on this connection. One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
$\widetilde{\mathcal{P}}(\mathbf{W}_{1},\mathbf{W}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{W}_{1},\beta\mathbf{W}_{2})$
Furthermore, this convergence is uniform as a function of the vectors $\mathbf{W}_{1},\mathbf{W}_{2}$. See Lemma A.1 for details.
@@
For real-valued weights $\mathbf{W}=\bigl{(}W_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr{)}$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.7)-(1.8):
$G^{\mathbf{W}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}W_{(r,\ell)}.$
For $k\in\mathbb{Z}$, define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.9), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
@@
The same invariance in Theorem 1.4 holds when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$, and the maps $\mathcal{P}_{k_{j}}$ are replaced by $\widetilde{\mathcal{P}}_{k_{j}}$.
@@
The same braid relations in Theorem 1.6 hold for the maps $\widetilde{\mathcal{P}}_{k}$.
@@
The analogues of Theorems 1.7 and 1.8 hold, when $Z^{\mathbf{W}}$ is replaced by $G^{\mathbf{W}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $W_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $W_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.13-1.14 below).
The details of the proofs of these theorems are given in Appendix B.

### v3

@@ 1.3. Invariance of periodic polymer partition functions
Our next main result gives the invariance of the polymer partition functions under the iterated Pitman transform defined in (1.2). For $\sigma\in\mathcal{S}_{\mathbb{Z}}$, we let $\mathcal{W}^{\sigma}$ be the set of tuples $(U,V)\in\Psi$, such that
[…]
See Figure 4 for an example.
Figure 4. An example of an pair $(U,V)\in\mathcal{W}^{\sigma}$, where $U$ and $V$ each contain 33 elements. Each column corresponds to a periodic vector of weights $\mathbf{X}_{k}$. The gray horizontal lines are used to indicate a periodic environment with $N=4$. In this figure, the column parameters of columns $3,5$, and 66 (blue) are permuted, and the column parameters in columns 88 and 99 (red) are permuted. The condition is that, for each of the three end points $(x,y)$ in $U$ (lower left), $\sigma(\mathbb{Z}_{<x})=\mathbb{Z}_{<x}$, while for each of the three end points $(x,y)$ in $V$ (upper right), $\sigma(\mathbb{Z}_{>x})=\mathbb{Z}_{>x}$. The multi-path partition function for this pair is preserved after changing the weights from $\mathbf{X}$ to $\sigma\mathbf{X}$, and the distributional equality in Theorem 1.5 holds jointly over all end point pairs $(U,V)\in\Psi$ satisfying this property.
@@
Let $N\in\mathbb{N}$, and let $\mathbf{X}=(\mathbf{X}_{k}\in\mathbb{R}^{\mathbb{Z}_{N}}\mid k\in\mathbb{Z})=(X_{k,\ell}\mid k\in\mathbb{Z},\ell\in\mathbb{Z}_{N})$ be a collection of real-valued weights corresponding to the columns of the lattice. Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$ by the condition that $X_{k,j}=X_{k,\ell}$ whenever $j\equiv\ell\mod N$. Then, if $\sigma\in\mathcal{S}_{\mathbb{Z}}$ and $(U,V)\in\mathcal{W}^{\sigma}$,
$Z^{\mathcal{P}_{\sigma}\mathbf{X}}(V\mid U)=Z^{\mathbf{X}}(V\mid U).$ (1.11)
@@
Theorem 1.4 is proved in Section 3. It is the first of its kind for a polymer in a periodic environment. The invariance under Pitman transformations in other contexts has been widely studied. The concept can be traced back to queuing theory from the works [WEB79, TW87], where it was shown that the law of the departures process from two $\cdot/M/1$ queues in series is invariant under swapping of the service rates of the two queues. In the polymer context, Corwin [COR02] used the work of Noumi and Yamada [NY04] to prove an analogous result to our Theorem 1.4 for a full-line version of the discrete Pitman transform. The proof of Theorem 1.4 likewise uses the machinery from [NY04], but we instead work with $\mathbb{Z}\times\mathbb{Z}$ periodic matrices instead of finite-dimensional matrices. The matrix relations in Proposition 2.7 that are key to proving Theorem 1.1(ii) are also the key to the proof of Theorem 1.4.
@@
Theorem 1.4 limits to a zero-temperature analogue in Theorem 1.9. This zero-temperature limit is the periodic analogue of invariances shown in [DOV22b, DNV22a, DV24, BEM+25]. The result in [DOV22b] is particularly significant, as it was key in constructing the directed landscape–the central limiting object of the KPZ universality class.
@@
It would be interesting to see if Theorem 1.4 could be the starting point for constructing and proving convergence to the periodic directed landscape. However, we note that the action of this transform is qualitatively different than that in [DOV22b]. Specifically, in the full-line case, the action of the iterated Pitman transforms and independent collection of Brownian motions and transforms them to the Brownian melon–a set of nonintersecting paths whose geodesic structure is more amenable to asymptotic analysis. By contrast, in this periodic setting when we take the weights to be i.i.d. with the inverse-gamma distribution, the transformed weights are equal in distribution to the original weights. This is the Burke property shown in [CGS26, Proposition 5.12], and extended to an inhomogeneous setting in Proposition 4.1. As a corollary, we obtain a permutation invariance result for the multi-path inhomogeneous inverse-gamma polymer below in Theorem 1.5. This extends the previous result of [BEM+25, Theorem 3.1] to a positive-temperature and multi-path setting. In this sense, the periodic Pitman transform is more analogous to a limiting case of the full-line Pitman transform, where the initial point of a random walk is sent to -$\infty$ and the output of the transform is recentered at 0 (see [DNV22a, DV24]). This limiting full-line setting is where we see analogous Burke properties, as in [OY01, OY02].
1.4. Permutation invariance
In this section, we consider an inhomogeneous, periodic inverse-gamma polymer. Let $\lambda>0$, and let and $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ be sequences of real numbers such that $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$. We take mutually independent weights $\mathbf{X}^{\mathbf{a},\mathbf{b}}=\bigl(X_{(i,j)}\mid i\in\mathbb{Z},\ j\in\mathbb{Z}_{N}\bigr)$ such that $X_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$ (see Definition 1.13 below). Extend these periodically to weights on $\mathbb{Z}\times\mathbb{Z}$.
@@ 1.4. Permutation invariance
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}_{N}}$ satisfying $a_{i}+b_{j}>0$ for $i\in\mathbb{Z}$ and $j\in\mathbb{Z}_{N}$, define periodic weights $\mathbf{X}^{\mathbf{a},\mathbf{b}}$ as above and extend them to $\mathbb{Z}\times\mathbb{Z}$. Then, for any finite permutation $\sigma\in\mathcal{S}_{\mathbb{Z}}$, we have
$\Bigl(Z^{\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma}\Bigr)\overset{d}{=}\Bigl(Z^{\mathbf{X}^{\sigma\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma}\Bigr),$
where $\sigma\mathbf{a}=(a_{\sigma(i)})_{i\in\mathbb{Z}}$.
The condition (1.10) ensures that, between the pairs of points in $(U,V)$, the column rates are either permuted among themselves or are left unchanged. See Figure 4.
We can also permute the $\mathbf{b}$ parameters as well, with doubly periodic weights $(X_{(i,j)}\mid i\in\mathbb{Z}_{M},j\in\mathbb{Z}_{N})$ for $M,N\in\mathbb{N}$, when we restrict to points that lie inside the box $\llbracket 0,M\rrbracket\times\llbracket 0,N\rrbracket$. In this case, the model isn’t actually periodic at all; it is simply a restriction of the full-space polymer to a finite grid. For finite permutations $\sigma,\tau\in\mathcal{S}_{\mathbb{Z}}$, define the set $\mathcal{W}^{\sigma,\tau}$ to be the set of tuples $(U,V)$, where $U,V$ are each sets of end point pairs in $\mathbb{Z}\times\mathbb{Z}$ satisfying the condition (1.10) for $\sigma$ as well as the condition
@@
The following is an extension of the recent result [BEM+25, Theorem 3.1] to multi-paths and to the positive-temperature case. A sketch of the proof for the extension to multi-paths in zero-+temperature was previously discussed in [BEM+25, Remark 4.3]. The technique follows the same outline as developed in [BEM+25], as we use a coupling of the weights given by applying a Pitman transform to neighboring rows and columns. However, the coupling we use in this proof is different because we treat the weights as periodic outside a large box and apply the periodic Pitman transform. The algebra involved in this positive-temperature case is also different.
@@
For vectors $\mathbf{a}\in\mathbb{R}^{\mathbb{Z}}$ and $\mathbf{b}\in\mathbb{R}^{\mathbb{Z}}$ satisfying $a_{i}+b_{j}>0$ for $i,j\in\mathbb{Z}$ define independent weights $\mathbf{X}^{\mathbf{a},\mathbf{b}}=\bigl(X_{(i,j)}:i,j\in\mathbb{Z}\bigr)$ such that $X_{(i,j)}\sim\text{Log-Inv-Gamma}(a_{i}+b_{j},\lambda)$ for all $i,j$. Then, for any finite permutations $\sigma,\tau\in\mathcal{S}_{\mathbb{Z}}$, we have
$\Bigl(Z^{\mathbf{X}^{\mathbf{a},\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma,\tau}\Bigr)\overset{d}{=}\Bigl(Z^{\mathbf{X}^{\sigma\mathbf{a},\tau\mathbf{b}}}(V\mid U):(U,V)\in\mathcal{W}^{\sigma,\tau}\Bigr),$
@@
Theorems 1.5 and 1.6 are proved in Section 4. When the initial point is fixed, invariances of this flavor can be deduced from the symmetry of Schur functions and Whittaker functions (see, for example, [BP08, COS+14, OSZ14, BC14]). A similar type of invariance for last-passage percolation was discovered in the work of Dauvergne [DAU22, Theorem 1.5], motivated by a shift-invariance result previously proved by Borodin, Gorin, and Wheeler [BGW22]. Connections between these results are discussed further in [BEM+25, Section 3].
@@ 1.5. Zero-temperature limits
We define the zero-temperature discrete periodic Pitman transform as the map $\widetilde{\mathcal{P}}:\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}\to\mathbb{R}^{\mathbb{Z}_{N}}\times\mathbb{R}^{\mathbb{Z}_{N}}$ defined by
$\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=\bigl(\widetilde{T}(\mathbf{X}_{1},\mathbf{X}_{2}),\widetilde{D}(\mathbf{X}_{1},\mathbf{X}_{2})\bigr),\quad\text{where}$ (1.13)
$\widetilde{T}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{2,i-1}+\max_{j\in\mathbb{Z}_{N}}Y_{[i-1,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{[i,j]},\quad\text{and}$
$\widetilde{D}_{i}(\mathbf{X}_{1},\mathbf{X}_{2})$ $=X_{1,i+1}+\max_{j\in\mathbb{Z}_{N}}Y_{(i,j]}-\max_{j\in\mathbb{Z}_{N}}Y_{(i-1,j]},\quad\text{where}\quad Y_{\ell}=X_{1,\ell+1}-X_{2,\ell}.$
The mapping $\widetilde{D}$ is similar but not equivalent to a mapping used to describe the invariant measures for multi-type TASEP by Ferrari and Martin [FM07]. See [CGS26, Appendix F] for more on this connection. One immediately sees that we can obtain $\widetilde{\mathcal{P}}$ as a limiting case of the positive-temperature discrete periodic Pitman transform:
$\widetilde{\mathcal{P}}(\mathbf{X}_{1},\mathbf{X}_{2})=\lim_{\beta\to\infty}\frac{1}{\beta}\mathcal{P}(\beta\mathbf{X}_{1},\beta\mathbf{X}_{2}).$
@@
For real-valued weights $\mathbf{X}=\bigl(X_{(i,j)}\in\mathbb{R}\mid i,j\in\mathbb{Z}\bigr)$ corresponding to the vertices of the directed lattice $\mathbb{Z}^{2}$, we define the last-passage percolation (LPP) time for both a single pair of points and multi-paths analogously as in (1.8)-(1.9):
$G^{\mathbf{X}}(V\mid U)=\max_{\pi\in\Pi_{U,V}}\sum_{(r,\ell)\in\pi}X_{(r,\ell)}.$
For $k\in\mathbb{Z}$, we can also define the map $\widetilde{\mathcal{P}}_{k}:(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}\to(\mathbb{R}^{\mathbb{Z}_{N}})^{\mathbb{Z}}$ by (1.3), with $T,D$ replaced by $\widetilde{T},\widetilde{D}$.
@@
The same relations in Theorem 1.1 hold for the maps $\widetilde{\mathcal{P}}_{k}$.
Theorem 1.8 similarly allows us to define maps $\widetilde{\mathcal{P}}_{\sigma}$ for a general finite permutation $\sigma\in\mathcal{S}_{\mathbb{Z}}$.
@@
The same invariance in Theorem 1.4 holds when $Z^{\mathbf{X}}$ is replaced by $G^{\mathbf{X}}$, and the maps $\mathcal{P}_{\sigma}$ are replaced by $\widetilde{\mathcal{P}}_{\sigma}$.
@@
The analogues of Theorems 1.5 and 1.6 hold, when $Z^{\mathbf{X}}$ is replaced by $G^{\mathbf{X}}$ and, instead of log-inverse-gamma weights, the weights satisfy either $X_{(i,j)}\sim\text{Geom}(q_{i}r_{j})$ for all $i,j$, where $q_{i}r_{j}\in(0,1)$, or $X_{(i,j)}\sim\text{Exp}(a_{i}+b_{j})$ for all $i,j$, where $a_{i}+b_{j}>0$ (see Definitions 1.11-1.12 below).
The details of the limit transitions for the proofs of these theorems are given in Appendix A.2.